OpenStudy (1018):

how to find the lateral surface area of a hexagonal pyramid whose edges measures 20cm and the radius of a circle inscribed in the circle is 9sqrt3 cm.

OpenStudy (anonymous):

Hexagonal Pyramid Formula : Area of Base(A) = (6/2)as = 3as Surface Area of Pyramid = 3as + 3sl = A + 3sl Volume of Pyramid = abh where a = apothem length, s,b = side, h = height and l = slant height